OpenStudy (anonymous):

A particular form of electromagnetic radiation has a frequency of 5.42x10^15 Hz. What is the wavelength is nanometers? meters?

OpenStudy (xishem):

\[c = \lambda \nu \rightarrow \lambda = \frac{c}{\nu}\]\[\lambda = \frac{3.00*10^8m*s^{-1}}{5.42*10^{15}s^{-1}}=5.54*10^8m\]And in nanometers...\[5.5*10^8m*(\frac{1*10^9nm}{1m})=55.4nm\]